# Fate of transient order parameter domain walls in ultrafast experiments

### v1

@@ Abstract
In ultrafast experiments, an optical pump pulse often generates transient domain walls of the order parameter in materials with spontaneous symmetry breaking, due to either a finite penetration depth of the light on a three-dimensional (3D) material, or a finite spot size on a two-dimensional (2D) material. We clarify the decaying process of such a domain wall that is caused by fluctuations of the order parameters. We study a generic system with $U(1)$-symmetric order, and those with an additional weak $Z_{2}$ ($U(1)$-symmetry-breaking) term, representing the charge-density-wave (CDW) orders in recent experiments. The decay process comprises two non-trivial stages. During the first stage, exponentially growing thermal fluctuations convert the domain wall into an interface with randomly distributed topological defects. In the second stage, the topological defects undergo a coarsening dynamics within the interface. For a 2D interface in the 3D system, the coarsening dynamics leads to a diffusive growth of the correlation length. For a one-dimensional (1D) interface in the 2D system with the weak $Z_{2}$ term, the correlation-length growth shows a crossover from diffusive to sub-diffusive behavior. Our theory provides a fundamental physical picture for the dynamics of pump-induced domain walls in ultrafast experiments.
@@
Introduction— Transient manipulation of order parameters by an ultrafast laser pulse has been one of the major topics in condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Recent experiments [25, 26, 27, 28] reported that in several CDW materials, a laser pulse (pump) created domain walls of the CDW order parameter. Thereby, the pump pulse heats the surface of the system, and the CDW order parameter $\psi$ in the surface region is suppressed in time from its initial equilibrium value $\psi_{\text{eq}}$ to zero, following a classical equation of motion. From the principle of inertia, $\psi$ may cross zero, and stabilize in the energy minimum at $-\psi_{\text{eq}}$ after the system is cooled down [25, 26, 27, 28], getting ‘inverted’. Due to the finite penetration depth of the pump pulse, only the order parameter in the surface region gets inverted, yielding a domain wall (white plane in the left panel of Fig. 1) between the surface region and the bulk at time $t=t_{0}$, see Sec. I of Ref. [29]. Within a mean-field theory of the time-dependent Ginzburg-Landau (TDGL) dynamics that neglects fluctuations, the domain wall undergoes no further evolution [25, 26, 28, 27].
@@
In this letter, we propose that due to fluctuations, the domain wall will decay following a two-stage dynamics. In the first stage, the domain wall decays by the exponential growth of unstable fluctuations [30, 16] of $\psi$ on the wall, which happens both in the $U(1)$ symmetric case, and in the system with a weak $Z_{2}$ term. The dynamics is characterized by increasing magnitude and correlation length of the fluctuations. At the end of the first stage, around a timescale $t\simeq t_{\rm c}$, the wall is transformed into an interface containing randomly distributed topological defects, as shown by the vortex strings in Fig. 1 for a 3D system. In 2D systems, the interface is 1D and the defects are vortices and antivortices.
@@
$U(1)$-symmetric model— We consider the “model A” dynamics [38, 39], whose mean-field version is also called the ‘TDGL equation’, together with a $U(1)$-invariant free energy:
@@
Here $\psi(\bm{r},t)=\psi_{1}+i\psi_{2}$ is the dimensionless complex order parameter, $\gamma$ is the relaxation rate, $E_{\text{c}}$ is the condensation energy density, $D$ is the space dimension, $\alpha\sim(T-T_{\text{c}})/T_{\text{c}}$ is negative at a temperature $T$ below the critical temperature $T_{\text{c}}$, and $\xi_{0}$ is the bare coherence length. The thermal noise field $\eta=\eta_{1}+i\eta_{2}$ arises from a thermal bath made of the other degrees of freedom, including high-energy electrons and phonons. We characterize the noise by a correlator $\langle\eta_{i}(\bm{r},t)\eta_{i^{\prime}}(\bm{r^{\prime}},t^{\prime})\rangle=2T(\gamma E_{\text{c}})^{-1}\delta_{ii^{\prime}}\delta(\bm{r}-\bm{r^{\prime}})\delta(t-t^{\prime})$ with an effective temperature $T$ of the bath and $i/i^{\prime}=1,2$. Here the Ginzburg parameter $\zeta=|\alpha|^{D/2-2}T/(E_{\text{c}}\xi_{0}^{D})$ plays the role of a dimensionless measure of the fluctuations in the non-equilibrium process as well as the equilibrium thermal fluctuations  [16, 40]. Eqs. (1)(2) effectively describe the dynamics of the mean-field order parameter and its fluctuations at temperatures close to $T_{\text{c}}$ [38, 40], in incommensurate CDW [14, 15], spin density waves (SDW) [41], excitonic order, and superconductivity [39, 42, 40, 16].
@@
where $a=1,2$, $\hat{L}_{1}=-\xi_{0}^{2}\partial_{z}^{2}+2\alpha+6\psi_{0}^{2}$, $\hat{L}_{2}=-\xi_{0}^{2}\partial_{z}^{2}+2\alpha+2\psi_{0}^{2}$, and $\nabla_{\bm{R}}^{2}=\partial_{x}^{2}+\partial_{y}^{2}$ is the Laplacian in $D-1$ dimensions. The eigenfunction of $\hat{L}_{a}+\xi_{0}^{2}\nabla_{\bm{R}}^{2}$ can be written as $\phi_{a}^{\bm{K}}(\bm{r})=u_{a}(z)e^{i\bm{K}\cdot\bm{R}}$ with eigenvalue $\epsilon_{a}+\xi_{0}^{2}K^{2}$, where $u_{a}(z)$ is the eigenfunction of $\hat{L}_{a}$ with eigenvalue $\epsilon_{a}$. Here $\bm{R}$($\bm{K}$) denotes the spatial coordinate (momentum) in the $(D-1)$-dimensional plane normal to $z$. Eigenvalues of $\hat{L}_{1}$ are all non-negative, while $\hat{L}_{2}$ has one negative eigenvalue, $\epsilon_{2}=\alpha$, corresponding to a bound-state eigen-function $u_{2}(z)=\text{sech}(z/\xi)$, see Sec. II A of Ref. [29]. It implies that the long-wavelength fluctuations in the imaginary part direction, decomposed as $\phi_{2}^{\bm{K}}(\bm{r})=\text{sech}(z/\xi)e^{i\bm{K}\cdot\bm{R}}$ with $K<\sqrt{|\alpha|}/\xi_{0}$, will grow exponentially in time with the exponent $\left(|\alpha|-\xi_{0}^{2}K^{2}\right)\gamma t$ [16].
With $\phi_{2}(\bm{r})=\varphi(\bm{R},t)\text{sech}(z/\xi)$, the correlation function $C(\bm{R},t)=\langle\varphi(\bm{0},t)\varphi(\bm{R},t)\rangle$ depicts the real space structure of the fluctuation within the interface. From Eq. (3), it has the following analytical expression for $\gamma t\gg 1/|\alpha|$ (see Sec. II B of Ref. [29]):
@@
with the modified Bessel function $I_{0}$ of the first kind. Therefore, the fluctuation magnitude grows exponentially with the correlation length growing as $L(t)=4\xi_{0}\sqrt{\gamma t}$, a diffusive behavior. Note that due to the initial fluctuations $\langle\phi_{2}(\bm{r})\phi_{2}(\bm{r^{\prime}})\rangle\sim 1/|\bm{r}-\bm{r^{\prime}}|$ at long distances, the spatial correlation crossovers from an exponential decay at $R\ll L(t)$ to a power law decay $C\left(\bm{R},t\right)\propto L(t)/R$ at $R\gg L(t)$. The first-stage dynamics ends at a crossover time $t_{\text{c}}$ when the fluctuation becomes so large that the linear approximation in Eq. (3) fails, and the $|\psi|^{4}$ term in Eq. (2) starts to stop its growth. One may estimate this timescale by $\langle|\psi(\bm{r},t_{\text{c}})|^{2}\rangle\sim|\alpha|$, which gives $C(\bm{0},t_{\text{c}})\sim|\alpha|$ and $\gamma t_{\text{c}}\sim(2|\alpha|)^{-1}\ln(\zeta^{-1})$. Note that $\gamma t_{\text{c}}$ thus obtained is much larger than $1/|\alpha|$ given that $\zeta\ll 1$.
@@
At $t\gtrsim t_{\text{c}}$, the second stage of dynamics takes over, which has been argued to be a coarsening dynamics [30] of the vortex strings. Thereby, distributions of the vortex strings at different times show a statistical self-similarity, which is characterized by a growing correlation length $L(t)$, or the typical inter-string distance. The coarsening dynamics with $L(t)\propto t^{p}$ and $p>1/2$, $p=1/2$, and $p<1/2$ is referred to as superdiffusive, diffusive, and subdiffusive dynamics, respectively. Eq. (1) implies the equation of motion for the length scale:
@@
see Refs. [30, 32, 31, 31, 34] and Sec. III B of Ref. [29]. Here $\lambda(L)\frac{dL}{dt}$ and $F_{\text{d}}(L)$ are the frictional force with a friction coefficient $\lambda(L)$ and the driving force acting on a unit length of the vortex string, respectively. Eq. (5) is nothing but a balance equation between the two forces. Consider a circular vortex-string loop with the diameter $L$, a typical topological defect, its energy is $E\sim|\alpha|E_{\text{c}}\xi_{0}^{2}L\ln(L/\xi)$ [30, 29]. Therefore, the driving force that contracts the loop is $F_{\text{d}}(L)\sim L^{-1}dE/dL\sim|\alpha|E_{\text{c}}\xi_{0}^{2}L^{-1}\ln(L/\xi)$ in the large $L$ limit. The friction coefficient $\lambda(L)\sim\gamma^{-1}E_{\text{c}}|\alpha|\ln(L/\xi)$ is derived from the damping term [30, 29], i.e. first-order time derivative term in Eq. (1). Combining these two, we obtain a solution of Eq. (5) in the large-$L$ limit as $L(t)\sim\xi_{0}\sqrt{\gamma t}$, which manifests a diffusive coarsening dynamics.
@@
We numerically simulated Eq. (1) with the domain-wall initial condition and observed the coarsening dynamics. To evaluate the length scale $L(t)$ in the simulation, we count a number $N$ of vortices in the $y$-$z$ planes for each discrete $x$, average $N$ over different $x$, and define $L=L_{y}/\langle N\rangle$, where $L_{y}$ is the system size along $y$. Fig. 3(a) shows $L$ versus time by a log-log plot in the interval of $\log_{10}(L/\xi)\in(1,1.2)$, from which we estimate the exponent $p$ by a linear fit $\log_{10}\left(L/\xi\right)=p\log_{10}(|\alpha|\gamma t)+c_{0}$. The estimate gives $p=0.45$ for the 2D interface of a 3D system. The discrepancy from the theoretical result may be because the simulation has not entered the long-time region.
One can define a structure factor $S(\bm{K},t)$ on the interface by a Fourier transform of $C(\bm{R},t)$ [36, 37], which may be experimentally measurable by time-resolved X-ray diffraction [43, 37]. The correlation length $L(t)$ could be fitted from the dynamical scaling form $S(\bm{K},t)=g(KL)$  [30].
@@
Figure 3: Numerically computed $\log_{10}\left(L/\xi\right)$ as a function of $\log_{10}(|\alpha|\gamma t)$ for $0.00012\zeta=$ 0.0001, and the linear fitting. In the numerical simulation, space and time are discretized by step sizes $0.5\xi$ and $0.03(|\alpha|\gamma)^{-1}$. Data points are selected within $\log_{10}(L/\xi)\in(1,1.2)$. (a) The 3D $U(1)$ symmetric case. (b) The 3D $U(1)$ breaking case with $\beta/|\alpha|=0.1$. Both (a) and (b) are simulated in a $L_{x}\times L_{y}\times L_{z}$ grid with $L_{z}/\xi=60$, $L_{x}/\xi=L_{y}/\xi=1280$. (c)(d) The 2D case with the $Z_{2}$ term $\beta/|\alpha|$ being $0.001$ (c) and $0.1$ (d). The phase distribution of a topological defect is sketched in the inset of (d) with the same colorbar as Fig. 2(b,c). Both (c) and (d) are simulated in a $L_{y}\times L_{z}$ grid with $L_{z}/\xi=120$, $L_{y}/\xi=38400$.
@@
The effects of a $Z_{2}$ term— To model the case without the $U(1)$ symmetry, we add a term $2\beta\psi_{2}^{2}$ (‘$Z_{2}$ term’) with $\beta>0$ to the free energy density in Eq. (2), reducing the $U(1)$ symmetry to a $Z_{2}\otimes Z_{2}$ symmetry under the transformation $\psi_{1(2)}\rightarrow\pm\psi_{1(2)}$. Due to the $Z_{2}$ term, the low-temperature equilibrium state is long-range ordered even for a 2D system so that a well-defined 1D domain wall could be generated. This model describes commensurate CDW/SDW [44, 45] and excitonic insulators [46], etc. As shown below, the decay process of the domain wall for the small $\beta$ case is characterized by a similar two-stage dynamics as the $U(1)$ symmetric case, while the scaling behavior may be different.
@@
For the first-stage dynamics, the $Z_{2}$ term modifies $\hat{L}_{2}$ in Eq. (3), $\hat{L}_{2}=-\xi_{0}^{2}\partial_{z}^{2}+2(\alpha+\beta)+2\psi_{0}^{2}$, so that the lowest eigenvalue of $\hat{L}_{2}$ becomes $\alpha+2\beta$. For $\beta>|\alpha|/2$, the eigenvalue is positive, and the domain wall is stable against any fluctuation of the imaginary part of $\psi$. For $\beta<|\alpha|/2$, the mode $\phi_{2}^{\bm{K}}(\bm{r})=\text{sech}(z/\xi)e^{i\bm{K}\cdot\bm{R}}$ with $K<\sqrt{|\alpha|-2\beta}/\xi_{0}$ grows exponentially in time. The growth is characterized by the correlation function $C(\bm{R},t)$. The correlation function in the long-time regime with $\gamma t\gg 1/\beta$ and $\gamma t\gg 1/(|\alpha|-2\beta)$ is computed from Eq. (3) as [29]
@@
where $c_{1}\approx\sqrt{|\alpha|/\beta}$ if $\beta\ll|\alpha|$. Similar to Eq. (4), one has a growing correlation length $L(t)\sim\xi_{0}\sqrt{8\gamma t}$. The crossover time to the second-stage dynamics is $\gamma t_{\text{c}}\sim(2|\alpha|)^{-1}\ln(\zeta^{-1})$ for $\beta\ll|\alpha|$, given that $\ln(\zeta^{-1})\gg|\alpha|/\beta$ such that $t_{\text{c}}$ is in the long-time regime.
The coarsening dynamics of the topological defects with the $Z_{2}$ term is also described by the balance equation Eq. (5), while the friction coefficient $\lambda(L)$ and the driving force $F_{\text{d}}(L)$ depends on a spatial profile of the topological defect that is deformed by the $Z_{2}$ term. In the case of $\beta\ll|\alpha|$, the free energy density $f$ is dominated by its phase part,
@@
Even with the small $\beta$, the topological defect is no longer a perfect circular vortex, but connects two different sine-Gordon solitons $\theta_{0}^{\pm}=\text{arg}\left(\sinh(\sqrt{2\beta}z/\xi_{0})\pm i\right)$  [29, 47], which is sketched in the inset of Fig. 3(d). When the distance $r$ from the vortex core in the 2D plane is much smaller than a length scale $\xi_{0}/\sqrt{2\beta}$, the spatial profile approaches the form of a circular symmetric vortex. For $r\gg\xi_{0}/\sqrt{2\beta}$, the profile $\theta_{0}^{\pm}$ approaches the sine-Gordon solitons as $\theta\approx\theta_{0}^{\pm}+s(\bm{r})$, where $s(\bm{r})\sim e^{-r\sqrt{2\beta}/\xi_{0}}$ decays exponentially.
As a result, the coarsening dynamics crossovers between two asymptotic time domains. In the early-time domain with $L(t)\ll\xi_{0}/\sqrt{2\beta}$, adjacent defects see each other as circular symmetric vortices. Therefore, the driving force and frictional force are the same as the $U(1)$ symmetric case. In 2D systems, they are $F_{\text{d}}(L)\sim E_{\text{c}}|\alpha|\xi_{0}^{2}/L$, and $\lambda(L)\sim E_{\text{c}}|\alpha|\ln(L/\xi)/\gamma$, respectively. From Eq. (5), one has $L(t)\sim\xi_{0}\sqrt{\gamma t/\ln(|\alpha|\gamma t)}$, which is nearly diffusive in the large-$\gamma t$ limit. In 3D systems, the calculation is explained below Eq. (5). In the late-time domain with $L(t)\gg\xi_{0}/\sqrt{2\beta}$, the adjacent vortices see each other as the sine-Gordon solitons. Thereby, the friction coefficient becomes a constant $\lambda\sim E_{\text{c}}|\alpha|/\gamma$. The driving force in the 2D system becomes short-range attraction between vortices and antivortices with $F_{\text{d}}(L)\sim\xi_{0}E_{\text{c}}|\alpha|\sqrt{2\beta}e^{-L\sqrt{2\beta}/\xi_{0}}$, while the driving force in the 3D system is dominated by the string tension with $F_{\text{d}}(L)\sim\xi_{0}^{2}E_{\text{c}}|\alpha|L^{-1}$, see Sec. III B of Ref. [29]. This results in the following scaling for the late-time domain of the coarsening dynamics,

### v2

@@ Abstract
In ultrafast experiments, an optical pump pulse often generates transient domain walls of the order parameter in materials with spontaneous symmetry breaking, due to either a finite penetration depth of light on a three-dimensional (3D) material, or a finite spot size on a two-dimensional (2D) material. We show that the domain wall decays due to unstable order parameter fluctuations. We study a generic system with $U(1)$-symmetric order, and those with an additional weak $Z_{2}$ ($U(1)$-symmetry-breaking) term, representing the charge-density-wave (CDW) orders in recent experiments. During the first stage of the decay dynamics, exponentially growing thermal fluctuations convert the domain wall into an interface with randomly distributed topological defects. In the second stage, the topological defects undergo a coarsening dynamics within the interface. For a 2D interface in a 3D system, the coarsening dynamics leads to a diffusive growth of the correlation length. For a one-dimensional (1D) interface in a 2D system with the weak $Z_{2}$ term, the correlation-length growth shows a crossover from diffusive to sub-diffusive behavior.
Introduction
@@
Transient manipulation of order parameters by an ultrafast laser pulse has been one of the major topics in condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Recent experiments [27, 28, 29, 30, 31] reported that in several CDW materials, a laser pulse (pump) created domain walls of the CDW order parameter. There the pump pulse heats the surface of the system so that the CDW order parameter $\psi$ in the surface region is suppressed in time from its initial equilibrium value $\psi_{\text{eq}}$ to zero, following a classical equation of motion. From the principle of inertia, $\psi$ may cross zero, and stabilize in the energy minimum at $-\psi_{\text{eq}}$ after the system is cooled down [27, 28, 29, 30], getting ‘inverted’. Due to the finite penetration depth of the pump pulse, only the order parameter in the surface region gets inverted, yielding a domain wall (the white plane in the left panel of Fig. 1) between the surface region and the bulk at time $t=t_{0}$, see the Supplementary Note 1. Within a mean-field theory of the time-dependent Ginzburg-Landau (TDGL) dynamics that neglects fluctuations, the domain wall undergoes no further evolution [27, 28, 30, 29].
@@
In this work, we propose that due to fluctuations, the domain wall will decay following a two-stage dynamics. In the first stage, the domain wall decays by the exponential growth of unstable fluctuations [32, 17] of $\psi$ on the wall, which happens both in the $U(1)$ symmetric case, and in the system with a weak $Z_{2}$ term. The dynamics are characterized by increasing magnitude and correlation length of the fluctuations. At the end of the first stage, around a timescale $t\simeq t_{\rm c}$, the wall is transformed into an interface containing randomly distributed topological defects, as shown by the vortex strings in Fig. 1 for a 3D system. In 2D systems, the interface is 1D and the defects are vortices and antivortices.
@@
Results
$U(1)$-symmetric model
We consider the “model A” dynamics [40, 41], whose mean-field version is also called the ‘TDGL equation’, together with a $U(1)$-invariant free energy:
@@ Results
Here $\psi(\bm{r},t)=\psi_{1}+i\psi_{2}$ is the dimensionless complex order parameter, $\gamma$ is the relaxation rate, $E_{\text{c}}$ is the condensation energy density, $D$ is the space dimension, $\alpha\sim(T-T_{\text{c}})/T_{\text{c}}$ is negative at a temperature $T$ below the critical temperature $T_{\text{c}}$, and $\xi_{0}$ is the bare coherence length. The thermal noise field $\eta=\eta_{1}+i\eta_{2}$ arises from a thermal bath made of the other degrees of freedom, including high-energy electrons and phonons. We characterize the noise by a correlator $\langle\eta_{i}(\bm{r},t)\eta_{i^{\prime}}(\bm{r^{\prime}},t^{\prime})\rangle=2T(\gamma E_{\text{c}})^{-1}\delta_{ii^{\prime}}\delta(\bm{r}-\bm{r^{\prime}})\delta(t-t^{\prime})$ with an effective temperature $T$ of the bath and $i/i^{\prime}=1,2$. Here the Ginzburg parameter $\zeta=|\alpha|^{D/2-2}T/(E_{\text{c}}\xi_{0}^{D})$ plays the role of a dimensionless measure of the fluctuations in the non-equilibrium process as well as the equilibrium thermal fluctuations  [17, 42]. Eqs. (1)(2) are reasonable approximations to the dynamics of the mean-field order parameter and its fluctuations at temperatures close to $T_{\text{c}}$ [40, 42] in incommensurate CDW [15, 16], spin density waves (SDW) [43], excitonic insulators (EI), and Bardeen–Cooper–Schrieffer (BCS) superconductors [41, 44, 42, 17]. A microscopic derivation of the free energy model from an electron-phonon coupling Hamiltonian is presented in Supplementary Note 7, which shows $E_{\rm c}\sim\nu T_{\rm c}^{2}$ with $\nu\sim p_{\rm F}^{D}/E_{\rm F}$ being the density of states of electrons contributing to the Fermi surface nesting. Here $p_{\rm F}$ and $E_{\rm F}$ denote the Fermi momentum and Fermi energy of the metallic state without the CDW formation. Note that the CDW often has different coherence lengths in different directions while Eq. (2) is written in a rescaled spatial coordinate so that it appears isotropic. Around $T_{\rm c}$, one has $\xi_{0}\sim v_{\rm F}/T_{\rm c}$ where $v_{\rm F}=E_{\rm F}/p_{\rm F}$ denotes the Fermi velocity [45, 46]. This leads to $\zeta\sim T_{\rm c}^{D-1}/E_{\rm F}^{D-1}|\alpha|^{D/2-2}$ [17] which is similar to SDW, EI, and BCS superconductors [43, 47, 48, 49].
@@
where $a=1,2$, $\hat{L}_{1}=-\xi_{0}^{2}\partial_{z}^{2}+2\alpha+6\psi_{0}^{2}$, $\hat{L}_{2}=-\xi_{0}^{2}\partial_{z}^{2}+2\alpha+2\psi_{0}^{2}$, and $\nabla_{\bm{R}}^{2}=\partial_{x}^{2}+\partial_{y}^{2}$ is the Laplacian in $D-1$ dimensions. The eigen-function of $\hat{L}_{a}-\xi_{0}^{2}\nabla_{\bm{R}}^{2}$ can be written as $\phi_{a}^{\bm{K}}(\bm{r})=u_{a}(z)e^{i\bm{K}\cdot\bm{R}}$ with eigenvalue $\epsilon_{a}+\xi_{0}^{2}K^{2}$, where $u_{a}(z)$ is the eigen-function of $\hat{L}_{a}$ with eigenvalue $\epsilon_{a}$. Here $\bm{R}$($\bm{K}$) denotes the spatial coordinate (momentum) in the $(D-1)$-dimensional plane normal to $z$. Eigenvalues of $\hat{L}_{1}$ are all non-negative, while $\hat{L}_{2}$ has one negative eigenvalue, $\epsilon_{2}=\alpha$, corresponding to a ‘bound-state’ eigen-function $u_{2}(z)=\text{sech}(z/\xi)$, see the Supplementary Note 2. It implies that the long-wavelength fluctuations in the imaginary part direction, written as $\phi_{2}^{\bm{K}}(\bm{r})\propto\text{sech}(z/\xi)e^{i\bm{K}\cdot\bm{R}}$ with $K<\sqrt{|\alpha|}/\xi_{0}$, will grow exponentially in time with the exponent $\left(|\alpha|-\xi_{0}^{2}K^{2}\right)\gamma t$ [17].
With $\phi_{2}(\bm{r})=\varphi(\bm{R},t)\text{sech}(z/\xi)$, the correlation function $C(\bm{R},t)=\langle\varphi(\bm{0},t)\varphi(\bm{R},t)\rangle$ depicts the real space structure of the fluctuation within the interface. From Eq. (3), it has the following analytical expression for $\gamma t\gg 1/|\alpha|$ (see the Supplementary Note 2):
@@
with the modified Bessel function $I_{0}$ of the first kind. Therefore, the fluctuation magnitude grows exponentially with the correlation length growing as $L(t)=4\xi_{0}\sqrt{\gamma t}$, a diffusive behavior. Note that due to the initial fluctuations $\langle\phi_{2}(\bm{r})\phi_{2}(\bm{r^{\prime}})\rangle\sim 1/|\bm{r}-\bm{r^{\prime}}|$ at long distances, the spatial correlation crosses over from an exponential decay at $R\ll L(t)$ to a power law decay $C\left(\bm{R},t\right)\propto L(t)/R$ at $R\gg L(t)$. The first-stage dynamics ends at a crossover time $t_{\text{c}}$ when the fluctuation becomes so large that the linear approximation in Eq. (3) fails, and the $|\psi|^{4}$ term in Eq. (2) starts to stop its growth. One may estimate this timescale by $\langle|\psi(\bm{r},t_{\text{c}})|^{2}\rangle\sim|\alpha|$, which gives $C(\bm{0},t_{\text{c}})\sim|\alpha|$ and $\gamma t_{\text{c}}\sim(2|\alpha|)^{-1}\ln(\zeta^{-1})$. Note that $\gamma t_{\text{c}}$ thus obtained is much larger than $1/|\alpha|$ given that $\zeta$ is small enough.
@@
At $t\gtrsim t_{\text{c}}$, the second stage of dynamics takes over, which has been argued to be a coarsening dynamics [32] of the vortex strings. Specifically, distributions of the vortex strings at different times show a statistical self-similarity, which is characterized by a growing correlation length $L(t)$, or the typical inter-string distance. The coarsening dynamics with $L(t)\propto t^{p}$ and $p>1/2$, $p=1/2$, and $p<1/2$ are referred to as superdiffusive, diffusive, and sub-diffusive dynamics, respectively. Eq. (1) implies the equation of motion for the length scale:
@@
see Refs. [32, 34, 33, 33, 36] and the Supplementary Note 3. Here $\lambda(L)\frac{dL}{dt}$ and $F_{\text{d}}(L)$ are the frictional force with a friction coefficient $\lambda(L)$ and the driving force acting on a unit length of the vortex string, respectively. Eq. (5) is nothing but a balance equation between the two forces. Consider a circular vortex-string loop with the diameter $L$, a typical topological defect, its energy is $E\sim|\alpha|E_{\text{c}}\xi_{0}^{2}L\ln(L/\xi)$ [32]. Therefore, the driving force that contracts the loop is $F_{\text{d}}(L)\sim L^{-1}dE/dL\sim|\alpha|E_{\text{c}}\xi_{0}^{2}L^{-1}\ln(L/\xi)$ in the large $L$ limit. The friction coefficient $\lambda(L)\sim\gamma^{-1}E_{\text{c}}|\alpha|\ln(L/\xi)$ is derived from the damping term [32], i.e. first-order time derivative term in Eq. (1). Combining these two, we obtain a solution of Eq. (5) in the large-$L$ limit as $L(t)\sim\xi_{0}\sqrt{\gamma t}$, which manifests a diffusive coarsening dynamics.
@@
We numerically simulated Eq. (1) with the domain-wall initial condition and observed the coarsening dynamics. To evaluate the length scale $L(t)$ in the simulation, we count a number $N$ of vortices in the $y$-$z$ planes for each discrete $x$, average $N$ over each $x$ in ten independent simulations, and define $L=L_{y}/\langle N\rangle$, where $L_{y}$ is the system size along $y$. Fig. 3(a) shows $L$ versus time by a log-log plot in the interval of $\log_{10}(L/\xi)\in(1,1.2)$, from which we estimate the exponent $p$ by a linear fit $\log_{10}\left(L/\xi\right)=p\log_{10}(|\alpha|\gamma t)+c_{0}$. The estimate gives $p=0.45$ for the 2D interface of a 3D system. The discrepancy from the theoretical result may be because the simulation has not entered the long-time region.
@@
Figure 3: Numerically computed $\log_{10}\left(L/\xi\right)$ as a function of $\log_{10}(|\alpha|\gamma t)$ for $0.00012\zeta=$ 0.0001, and the linear fitting. In the numerical simulation, space and time are discretized by step sizes $0.5\xi$ and $0.03(|\alpha|\gamma)^{-1}$. Data points are selected within $\log_{10}(L/\xi)\in(1,1.2)$. The results are statistical averages over ten independent simulations. (a) The 3D $U(1)$ symmetric case. (b) The 3D $U(1)$ breaking case with $\beta/|\alpha|=0.1$. Both (a) and (b) are simulated in a $L_{x}\times L_{y}\times L_{z}$ grid with $L_{z}/\xi=60$, $L_{x}/\xi=L_{y}/\xi=1280$. (c)(d) The 2D case with the $Z_{2}$ term $\beta/|\alpha|$ being $0.001$ (c) and $0.1$ (d). The phase distribution of a topological defect is sketched in the inset of (d) with the same colorbar as Fig. 2(b,c). Both (c) and (d) are simulated in a $L_{y}\times L_{z}$ grid with $L_{z}/\xi=120$, $L_{y}/\xi=38400$.
Effects of a $Z_{2}$ term
@@
To model the case without the $U(1)$ symmetry, we add a term $2\beta\psi_{2}^{2}$ (‘$Z_{2}$ term’) with $\beta>0$ to the free energy density in Eq. (2), reducing the $U(1)$ symmetry to a $Z_{2}\otimes Z_{2}$ symmetry under the transformation $\psi_{1(2)}\rightarrow\pm\psi_{1(2)}$. Due to the $Z_{2}$ term, the low-temperature equilibrium state is long-range ordered even for a 2D system so that a well-defined 1D domain wall could be generated. This model describes commensurate CDW/SDW [50, 46] and certain excitonic insulators [19], etc. As shown below, the decay process of the domain wall for the small $\beta$ case is characterized by a similar two-stage dynamics as the $U(1)$ symmetric case, while the scaling behavior may be different.
@@
For the first-stage dynamics, the $Z_{2}$ term modifies $\hat{L}_{2}$ in Eq. (3), $\hat{L}_{2}=-\xi_{0}^{2}\partial_{z}^{2}+2(\alpha+\beta)+2\psi_{0}^{2}$, so that the lowest eigenvalue of $\hat{L}_{2}$ becomes $\alpha+2\beta$. For $\beta>|\alpha|/2$, the eigenvalue is positive, and the domain wall is stable against any fluctuation of the imaginary part of $\psi$. For $\beta<|\alpha|/2$, the mode $\phi_{2}^{\bm{K}}(\bm{r})\propto\text{sech}(z/\xi)e^{i\bm{K}\cdot\bm{R}}$ with $K<\sqrt{|\alpha|-2\beta}/\xi_{0}$ grows exponentially in time. The growth is characterized by the correlation function $C(\bm{R},t)$. The correlation function in the long-time regime with $\gamma t\gg 1/\beta$ and $\gamma t\gg 1/(|\alpha|-2\beta)$ is computed from Eq. (3) as
@@
where $c_{1}\approx\sqrt{|\alpha|/\beta}$ if $\beta\ll|\alpha|$. Similar to Eq. (4), one has a growing correlation length $L(t)\sim\xi_{0}\sqrt{8\gamma t}$. The crossover time to the second-stage dynamics is $\gamma t_{\text{c}}\sim(2|\alpha|)^{-1}\ln(\zeta^{-1})$ for $\beta\ll|\alpha|$, given that $\ln(\zeta^{-1})\gg|\alpha|/\beta$ such that $t_{\text{c}}$ is in the long-time regime. Note that although the domain wall is locally stable for $\beta>|\alpha|/2$ in our calculation, it will finally decay in pump-probe experiments from other effects such as the nonzero curvature of the domain wall due to the finite spot size of lasers, see the Supplementary Note 8. It may also tunnel into another type of domain wall by nucleation processes, which we leave for future study.
The coarsening dynamics of the topological defects with the $Z_{2}$ term are also described by the balance equation Eq. (5), while the friction coefficient $\lambda(L)$ and the driving force $F_{\text{d}}(L)$ depends on a spatial profile of the topological defect that is deformed by the $Z_{2}$ term. In the case of $\beta\ll|\alpha|$, the free energy density $f$ is dominated by its phase part,
@@
Even with the small $\beta$, the topological defect is no longer a perfectly circular vortex, but connects two different sine-Gordon solitons $\theta_{0}^{\pm}=\text{arg}\left(\sinh(\sqrt{2\beta}z/\xi_{0})\pm i\right)$  [51], which is sketched in the inset of Fig. 3(d). When the distance $r$ from the vortex core in the 2D plane is much smaller than a length scale $\xi_{0}/\sqrt{2\beta}$, the spatial profile approaches the form of a circular symmetric vortex. For $r\gg\xi_{0}/\sqrt{2\beta}$, the profile approaches the sine-Gordon solitons as $\theta\approx\theta_{0}^{\pm}+s(\bm{r})$, where $s(\bm{r})\sim e^{-r\sqrt{2\beta}/\xi_{0}}$ decays exponentially.
As a result, the coarsening dynamics crosses over between two asymptotic time domains. In the early-time domain with $L(t)\ll\xi_{0}/\sqrt{2\beta}$, adjacent defects see each other as circular symmetric vortices. Therefore, the driving force and frictional force are the same as the $U(1)$ symmetric case. In 2D systems, they are $F_{\text{d}}(L)\sim E_{\text{c}}|\alpha|\xi_{0}^{2}/L$, and $\lambda(L)\sim E_{\text{c}}|\alpha|\ln(L/\xi)/\gamma$, respectively. From Eq. (5), one has $L(t)\sim\xi_{0}\sqrt{\gamma t/\ln(|\alpha|\gamma t)}$, which is nearly diffusive in the large-$\gamma t$ limit. In 3D systems, the calculation is explained below Eq. (5). In the late-time domain with $L(t)\gg\xi_{0}/\sqrt{2\beta}$, the adjacent vortices see each other as the sine-Gordon solitons. Thereby, the friction coefficient becomes a constant $\lambda\sim E_{\text{c}}|\alpha|/\gamma$. The driving force in the 2D system becomes short-range attraction between vortices and antivortices with $F_{\text{d}}(L)\sim\xi_{0}E_{\text{c}}|\alpha|\sqrt{2\beta}e^{-L\sqrt{2\beta}/\xi_{0}}$, while the driving force in the 3D system is dominated by the string tension with $F_{\text{d}}(L)\sim\xi_{0}^{2}E_{\text{c}}|\alpha|L^{-1}$, see the Supplementary Note 3. This results in the following scaling for the late-time domain of the coarsening dynamics,